# Conceptwm: A Diffusion Model Watermark for Concept Protection

### v1

@@ 4.5 Fidelity Preserving Perturbation Modulation
and $\|\sigma^{(i)}\|_{j}\leq\eta\quad\forall i\in\{1,..,N_{ft}\}$
where $x^{i}$ denotes the clean samples. Meanwhile, we need to ensure that the image quality and watermark loss are as minimal as possible. Our solution is to use watermark-free images for one-step DreamBooth to reduce the DreamBooth loss $\mathcal{L}_{\text{cond}}$ 1, while simultaneously applying PGD to enhance the adversarial loss $L_{ft}$. 2.
$\theta$ $\leftarrow\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$ (9)
$\theta^{\prime}$ $\leftarrow\theta.\text{clone}()$
$\varphi^{(i)}$ $\leftarrow\arg\max_{\delta}\mathcal{L}_{cond}(\theta^{\prime},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
Considering that the adversarial perturbation will affect the watermark, we also incorporate the decoded message loss as a constraint to optimize the perturbation.
$\mathcal{L}_{msg}=BCE(decoder(x+\sigma_{1}+\sigma_{2}^{(i)}),m_{target})$ (10)
@@
where $\sigma_{1}$ denotes the residual of the watermarked image and original image. The Fidelity Preserving Perturbation Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
@@
Figure 3: Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
Method Concept Oriented Concept Protection Low Data Requirement Fidelity Robustness
FID CLIP BitAcc BitAcc(Adv.) TPR TPR (Adv.)
None - - - 24.37 0.367 - - - -
Post-Processing
DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367
@@
Latent-based
Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
Merging Component
StableSig. $\times$ $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
Ours 24.58 0.365 95.68 89.32 0.994 0.884
Table 1: A comparison between our method and previous watermarking techniques is presented. The capacities for DWT-DCT-SVD, RivaGAN, StableSignature, and our method are 48 bits, 32 bits, 48 bits, and 48 bits, respectively. We control the false positive rate (FPR) at $10^{-6}$ and evaluate the true positive rate (TPR). Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case. ”Adv.” (Adversarial) refers to the average performance of images under various distortions.
5 Experiment
METHODS DISTORTIONS
Contrast Brightness BLUR Gauss Noise JPEG Crop&Rescale Vae.C Denoising Average
Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
@@ 5 Experiment
Ours 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
Dwtdctsvd 0.894 0.752 0.943 0.054 0.734 0.008 0.010 0.000 0.424
@@
Ours 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
Table 2: Our method’s performance against various types of attacks is compared with other watermarking schemes. The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression [5] and (8) Diffusion-based Reconstructive Attack [42].
Settings BitAcc(%) TPR
@@
Table 3: The table presents the results of our method under different experimental parameters. Our concept watermarking approach remains effective across various generation steps, different samplers, VAE versions, and generated shapes.
TRAIN TEST “a photo of <concept >person” “a portrait of <concept >person”
@@
Table 4: An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization. The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
@@
Table 5: The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
5.1 Implementation Details
Datasets. To evaluate the effectiveness of the proposed methods. For subject-driven generation, we select the face dataset CelebA-HQ [47]. For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period. Each image is resized and center-cropped to a resolution of 512 × 512.
Training Configurations. By default, we use the Stable Diffusion (v2.1) as the pretrained generator. Unless stated otherwise, the training instance prompt for subject-driven generation is ”a photo of <concept >person” and the prior prompt is ”a photo of person”. The training instance prompt for style-driven generation is ”a painting in the style of <concept >” and the prior prompt is ”a painting”. For PGD, we use $\alpha$ = 1e-5 and the default noise budget $\eta$ = 1e-3. The learning rate of the dreambooth is set 5e-6.
Evaluation Metrics. We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5. For traceability performance, we assess bit accuracy. To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) [13] and Structural Similarity Index (SSIM) [33] to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) [24] and Perceptual Image Quality Evaluator (PIQE) [32] for evaluating overall image quality.
Watermarking baselines. We select four typical baselines: two official watermark of Stable Diffusion [26] called DwtDctSvd [6], and RivaGAN [39],Vae-based method Stable Signature [9] and Latent-based method Tree-ring [34].
5.2 Watermark Performance Comparison
We evaluate our method against six typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, where the False Positive Rate (FPR) is set to 1%, and we assess the True Positive Rate (TPR) on 1,000 watermarked images. (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks. In this experiment, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing. Each user generates 10 images, creating a dataset of 10,000 watermarked images. During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images. The user with the highest Bit Accuracy is considered the traced user. The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks. In terms of Bit Accuracy, our method outperforms the best baseline by around 6.3%. This improvement is due to the widespread diffusion of the watermark throughout the latent space, creating a strong binding between the watermark and the image semantics.
5.3 Fidelity
Table 1 presents a comparison of our method with other baseline methods. For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) [12] for assessing image diversity and semantic relevance calculated on 5000 images from the COCO 2017 validation set [20] to assess image quality. Additionally, we also use CLIP score [25], a method for measuring image similarity, which provides results more aligned with human judgment.
Main Results. The performance of this method is relatively lower compared to other methods. It outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02). This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance. Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods.
5.4 Watermark Robustness against Attack
Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing. The main results are shown in Table 2.
Main Results. For each image processing attack, we report the average bit accuracy in Table 2. We see that our Conceptwm is indeed robust across all the transformations with the bit accuracy all above 0.9. Conceptwm remarkably outperforms existing multi-bit watermarks on the average performance with more than 6.3% boost compared with the state-of-the-art (SoTA) results. Since our method embeds the watermark into the entire latent space and trains the distortion layer, it has a clear advantage over contrastive watermarking schemes when facing various image processing attacks. Please refer to the Supplementary Materials for more results.
5.5 Concept Protection Evaluation
@@
Textual Inversion and DreamBooth with LoRA. Textual Inversion teaches new concepts by adjusting a word vector, rather than fine-tuning the entire model. LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency and is widely adopted within the community. The relevant results are shown in the Table 4. Our method shows better performance in mitigating the degradation effect on generation with LoRA. Since Textual Inversion does not inject concepts directly into the model, the reduction in quality is not significant and the semantic quality of the generation decrease.
@@
Inference Steps. Users may influence watermark extraction under different inference parameters. We generated images of specific concepts at different inference steps and performed watermark extraction. The results, shown in Table 3, indicate that the inference steps have little impact on watermark extraction.
@@
Sampling Method. We evaluate our watermark on DDIM [30], DPM-Solver multi-step [22], Euler and Heun samplers [16]. The results show that our watermark is minimally affected by the choice of sampler.
@@
6 Discussion
Limitations. First, our method exhibits limited detection performance under certain prompts. As shown in Figure 3, when the model generates poor results for specific concepts under certain prompts, the detection accuracy cannot be guaranteed. This issue requires more precise and improved personality techniques or larger datasets and prompts. Second, the adversarial perturbations generated by our proposed method have limited effectiveness against the latest purification methods [45, 2]. Since our approach involves embedding concepts, watermarks, and adversarial perturbations into the diffusion model, there are more constraints to consider, which makes our adversarial perturbations more limited compared to the state-of-the-art image adversarial perturbation methods [19, 31]. In the future, we will explore ways to better embed adversarial perturbations into the Diffusion model.
7 Conclusion
This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts. Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results. We have evaluated the stability and adversarial effectiveness of the watermarking method we designed, demonstrating that our defense remains effective under adverse conditions. In the future, our goal is to improve the quality and robustness of the watermark and consider more potential threats.

### v2

@@ 3.3. Adversarial Watermark Modulation
$\mathcal{L}_{ft}(\theta,x_{0})=\mathbb{E}_{x_{0},t}\|\epsilon-\epsilon_{\theta}(x_{t+1},t,c)\|_{2}^{2}+\lambda\|\epsilon^{\prime}-\epsilon_{\theta}(x_{t+1}^{\prime},t,c_{cls})\|_{2}^{2}$
where $\epsilon$ and $\epsilon^{\prime}$ are both sampled from the standard normal distribution $\mathcal{N}(0,\mathbf{I})$, The noisy variable $x_{t+1}^{\prime}$ is generated from the original stable diffusion model based on the prior prompt $c_{\text{cls}}$. $\lambda$ emphasizes the importance of the prior term in the generation process. The term $\lambda$ emphasizes the importance of the prior term in the generation process.
Algorithm 1 Adversarial Watermark Modulation
1:experiment AWM(($x_{0}$,$X_{B}$,$\theta$))
2:     for  $x_{b}\in X_{B}$ do
3:         c= “ a photo of $\langle$ concept $\rangle$ person ”
4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{b},c)$
5:     end for
6:     $s\leftarrow secret$
7:     $x_{w}=LatentWatermarkEmbedding(x_{0},s)$
8:     $\sigma_{1}=x_{w}-x_{0}$
9:     $Initialize$ $x_{0}^{0}\leftarrow x_{0}$
10:     for $t=1\ldots T$ do
11:         $\sigma_{2}^{*(i)}=\alpha\cdot sign(\nabla_{x}\mathcal{L}_{cond}(\theta,x_{0}^{t}+\sigma_{1}))$
12:         ${x_{0}}^{t+1}={x_{0}}^{t}+\sigma_{2}^{*(i)}$
13:     end for
14:     $\varphi=x_{0}^{T}+\sigma_{1}-x_{0}$
15:return $\varphi$
Adversarial Watermark Modulation. The goal of this modulation is to maximize $L_{cond}$ in Eq. (3) to hinder personalized finetuning and minimize $L_{ft}$ in Eq. (4) to learn the concept. Our solution is to use watermark-free images $\mathcal{X}_{A}$ for one-step DreamBooth training to reduce the DreamBooth loss $\mathcal{L}_{\text{ft}}$, while simultaneously applying Projected Gradient Descent (PGD) (Madry, 2017) to enhance the adversarial loss $L_{cond}$. The model used for PGD is trained with clean referenced images $\mathcal{X}_{B}$ in a one-step dreambooth training process, rather than the images used for concept learning. The purpose of this approach is to allow the model to learn adversarial perturbations without affecting the learning of the concept. The specific procedure is defined by Eq.s (5) and (6), where $\sigma_{1}$ denotes the residual of the watermarked image and original image. $x^{i}$ denotes the clean samples.
(5) $\theta^{*}=\arg\min_{\theta}\sum_{x\in\mathcal{X}_{A}}\mathcal{L}_{ft}(\theta,x^{i})$
(6) $\varphi^{(i)}=\arg\max_{\sigma_{2}^{(i)}}\mathcal{L}_{cond}(\theta^{*},x^{(i)}+\sigma_{1}+\sigma_{2}^{(i)})$
@@
The adversarial noise $\sigma^{*(i)}$ is within the constraint range of $\eta$. The Adversarial Watermark Modulation outputs an adversarial watermark $\varphi$ as the parameter for Concept-oriented Watermark Fine-tuning to facilitate learning.
Algorithm 2 Efficient Concept Watermark Finetuning
1:experiment $CWF$($X_{A}$,$X_{B}$,$\theta$,E)
2:     for  $x_{0}\in X_{A}$ do
3:         c=“ a photo of $\langle$ concept $\rangle$ person ”
4:         $\theta=\operatorname{TrainOneStepDreamBooth}(\theta,x_{0},c)$
5:         $\theta^{*}\leftarrow\text{ copy parameters from }\theta$
6:         $\varphi=AWM(x_{0},X_{B},\theta)$
7:         $t\sim\mathrm{Uniform}(\{1,\ldots,T\})$
8:         $z_{0}=E(x_{0})$
9:         $z_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
10:         $z_{t}^{\mathrm{w}}\leftarrow\sqrt{\bar{\alpha}_{t}}(z_{0}+\varphi)+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$
11:         Take gradient step on $\nabla_{\theta}\|{\epsilon_{\theta}}(z_{t}^{\mathrm{w}},t,c)-{\epsilon_{\theta^{*}}}(z_{t},t,c)\|^{2}$
12:     end for
13:return $\theta$
[…]
3.4. Efficient Concept Watermark Finetuning
We aim to integrate adversarial watermarking into concept learning in this module. The primary challenges are as follows: The availability of fine-tuned images for specific concepts is often limited, hindering the model’s ability to learn the distinct watermark patterns. The introduction of adversarial perturbations can interfere with the DreamBooth training process, as these protective perturbations prevent the model from effectively learning the concept. Our core idea is to decouple concept learning from adversarial watermark learning. In the DreamBooth stage, the U-Net learns the specific concept. In the adversarial watermark learning stage, the model is fed with both watermarked and non-watermarked images, and by using the difference in the predictions as a loss, it guides the generative model in learning the watermark embedding pattern.
The training objective of this stage is defined by Eq. (7), where $\varphi$ represents the adversarial perturbation containing the watermark generated through AWM, $\epsilon_{\theta}^{*}$ denotes the fixed model, and $\epsilon_{\theta}$ refers to the trained model.
(7) $\mathcal{L}_{\mathrm{cwft}}(\theta,x_{t})$ $=\mathbb{E}_{t,c,x}\Big{[}\|\boldsymbol{\epsilon}_{\theta}(x_{t}+\varphi,t,c)-\boldsymbol{\epsilon}_{\theta}^{*}(x_{t},t,c\big{)}\|^{2}\Big{]}$
Algorithm 2 outlines the finetuning procedure for the U-Net denoising model $\theta$ with three components: target concept images $X_{A}$, clean reference images $X_{B}$, and the VAE encoder $E$. The process has three phases: In concept specialization phase (Lines 1-5), The model $\theta$ is fine-tuned on $X_{A}$ to produce a specialized model $\theta^{*}$ that captures the target concept while maintaining general denoising capabilities. In adversarial watermark generation phase (Line 6), an adversarial perturbation is computed through the Fidelity-preserving Latent Watermark in section 3.2. In watermark finetuning phase (Lines 7-13), the model is fed with both watermarked and non-watermarked images, and the difference in the predictions is used as a loss to guide the generative model in learning the watermark embedding pattern.
[…]
@@ 3.4. Efficient Concept Watermark Finetuning
Figure 4. Demonstration of the generated images. Below each image, we display the corresponding Bit Accuracy, along with the image quality metrics PIQE and NIQE.
4. Experiment
4.1. Implementation Details
Datasets. To evaluate the effectiveness of the proposed methods, for subject-driven generation, we select the face dataset CelebA-HQ (Zhu et al., 2022). For style-driven generation, we select 12 artists with distinct styles from WikiArt  [^1] https://www.wikiart.org, with each artist contributing 10 artworks that are consistent in terms of both artistic style and time period. Each image is resized and center-cropped to a resolution of 512 × 512. In each experiment, a total of 5–8 carefully selected images were utilized for DreamBooth fine-tuning.
[…]
Type Method Concept Oriented Concept Protection Fidelity Robustness
FID $\uparrow$ CLIP $\uparrow$ BitAcc $\uparrow$ BitAcc(Adv.) $\uparrow$ TPR $\uparrow$ TPR (Adv.$\uparrow$
None - - 24.37 0.367 - - - -
Post-Processing DwtDctSvd $\times$ $\times$ 25.01 0.363 100.0 71.47 1.000 0.367
@@ 4.1. Implementation Details
Latent-based Tree-Ring $\times$ $\times$ 25.69 0.364 - - 1.000 0.821
VAE-based StableSig. $\times$ $\times$ 24.92 0.362 98.49 81.94 0.993 0.848
Concept-Oriented Ours(OLW-based) 24.06 0.357 98.68 90.03 0.998 0.884
Ours(FLW-based) 24.58 0.365 95.68 89.32 0.994 0.858
Table 1. A comparison between our method and previous watermarking techniques is presented. The capacities for DWT-DCT-SVD, RivaGAN, StableSignature, and our method are 48 bits, 32 bits, 48 bits, and 48 bits, respectively. We control the false positive rate (FPR) at $10^{-5}$ and evaluate the true positive rate (TPR). Since Tree-ring is a zero-bit watermark, bit accuracy cannot be calculated in this case. ”Adv.” (Adversarial) refers to the average performance of images under various distortions.
Distortion Layer during training. During training, several data augmentation techniques were employed to enhance the model’s robustness. We follow the setup of attack layers in the representative work Stable Signature (Fernandez et al., 2023). Specifically, JPEG compression was simulated by using the HiDDeN (Zhu et al., 2018) model’s JPEG distortion layer, with compression quality randomly chosen between 40 and 100 to simulate different levels of image compression. Random cropping was applied, resizing the cropped image within the range of [256, 512] pixels, and then scaling it back to 512×512 pixels to ensure variability in the region of interest. Gaussian blur was applied with kernel sizes randomly selected between [3, 9] and blur intensities ranging from 0 to 2, simulating distortions from defocus or motion blur. Gaussian noise with a mean of 0 and variance of 10 was added to simulate low-quality or noisy images. Color jittering adjusted brightness, contrast, and saturation within the range of [0.8, 1.25], and hue was randomly shifted between [-0.2, 0.2] to simulate lighting and color variations. These augmentations effectively improved the model’s adaptability to compression, noise, blur, and color variations, ensuring stable watermarking even under diverse image transformations.
Evaluation Metrics. We evaluate the detection performance using the true positive rate (TPR) at a false positive rate (FPR) of 10e-5. For traceability performance, we assess bit accuracy. To measure the quality of image generation, we compute the Peak Signal-to-Noise Ratio (PSNR) (Hore and Ziou, 2010) and Structural Similarity Index (SSIM) (Wang et al., 2004) to evaluate image distortion; and the Natural Image Quality Evaluator (NIQE) (Mittal et al., 2012) and Perceptual Image Quality Evaluator (PIQE) (Venkatanath et al., 2015) for evaluating overall image quality.
Watermarking baselines. We select four typical baselines: two official watermark of Stable Diffusion (Rombach et al., 2022) called DwtDctSvd (Cox et al., 2007), and RivaGAN (Zhang et al., 2019),VAE-based method Stable Signature (Fernandez et al., 2023) and Latent-based method Tree-ring (Wen et al., 2024). DwtDctSvd, RivaGAN, and Stable Signature all use 48-bit watermarks, while Tree-ring uses P-value as the evaluation method.
Metrics Methods Distortions
Contrast Brightness Blur Noise Jpeg Crop&Rescale Vae.C Denoising Average
Bit Accuracy(%$\uparrow$ Dwtdctsvd 88.41 86.80 92.42 65.87 83.32 54.02 55.74 49.53 72.01
@@
Ours (FLW-based) 94.26 93.46 93.68 92.42 93.63 90.26 90.12 84.63 91.34
Tpr(Fpr=$10^{-5}$$\uparrow$ Dwtdctsvd 0.894 0.752 0.943 0.054 0.734 0.008 0.010 0.000 0.424
@@
Ours (FLW-based) 0.968 0.951 0.947 0.958 0.949 0.912 0.910 0.734 0.916
Table 2. The performance of ConceptWM against various types of attacks is compared with other watermarking schemes. The attacks include (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020) and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023c).
4.2. Watermark Performance Comparison
We evaluate our method against 4 typical baselines on two tasks: (1) Detection: All methods are tested as single-bit watermarks with a unified watermark, with the False Positive Rate (FPR) is set to $10^{-5}$, and we assess the True Positive Rate (TPR) on 1,000 watermarked images. (2) Traceability: Except for the single-bit watermark Tree-Ring, all other methods are tested as multi-bit watermarks. Specifically, we simulate a scenario with 1,000 model users, each requiring a distinct watermark for model tracing. Each user generates 10 images, creating a dataset of 10,000 watermarked images. During testing, we calculate the Bit Accuracy by comparing the watermark of each user with the watermarked images generated by the user. The user with the highest Bit Accuracy is considered the traced user.
The comparison results, shown in Table 1, demonstrate that our watermarking method achieves superior robustness, significantly outperforming the baselines in both tasks. In terms of Bit Accuracy, our method outperforms the best baseline by around 6.3%. OLW exhibits strong performance in robustness experiments. However, it suffers from relatively unsatisfactory image quality (FID 24.06,CLIP 0.357). FLW demonstrates a relatively balanced result, maintaining good image quality while ensuring strong robustness. We found that there is a trade-off between image quality, robustness, and detection accuracy in watermark embedding. Due to extensive adversarial training on the watermark model, FLW achieves a bit accuracy of 95.68 in a non-attack environment, which is relatively lower compared to other watermarking schemes. To further improve bit accuracy in a non-attack setting, it would require considering either lower image quality or reduced robustness.
4.3. Fidelity
Table 1 presents a comparison of our method with other baseline methods. For fidelity evaluation metrics, we use the Fréchet Inception Distance (FID) (Heusel et al., 2017) for assessing image diversity and semantic relevance, which is calculated on 5000 images from the COCO 2017 validation set (Lin et al., 2014). Additionally, we use CLIP score (Radford et al., 2021), a method for measuring image similarity, which provides results more aligned with human judgment.
Settings BitAcc(%) TPR (Fpr=$10^{-6}$)
@@ 4.3. Fidelity
Table 3. The impact of inference settings on ConceptWM watermark detection.
Main Results. The performance of this method outperforms other methods in metrics like CLIP and FID, with minimal performance difference compared to the original model (FID difference of 0.21, CLIP difference of 0.02). This method is a localized fine-tuning of the model and does not significantly impact the model’s overall performance. Due to the smaller dataset and the more refined watermarking method, our method has a smaller impact on model generation quality compared to other methods. It can also be observed that, among the two variants of latent watermarking introduced, OLW and FLW, the image quality produced by OLW, with an FID of 24.06 and semantic quality (CLIP) of 0.357, is comparatively lower. In contrast, the image quality produced by FLW, has an (FID) of 24.58 and semantic quality (CLIP) of 0.365. This disparity arises because the watermark in FLW is contextually related to the image content, whereas the watermark in OLW is a fixed, image-independent watermark. Ultimately the latter leads to a degradation in image quality.
4.4. Watermark Robustness against Attack
Image Processing. We select eight representative types of image-level noise to evaluate the robustness of image processing: (1) Brightness, (2) Contrast, (3) Blur, (4) Gaussian Noise, (5) JPEG Compression, (6) Crop & Scale, (7) VAE Compression (Cheng et al., 2020), and (8) Diffusion-based Reconstructive Attack (Zhao et al., 2023d). These cover common image processing techniques, geometric transformations, and regeneration attacks encountered in real-world scenarios. The main results are shown in Table 2.
[…]
@@ 4.4. Watermark Robustness against Attack
Table 4. An experiment on generating images with adversarial watermarks and fine-tuning them with different model versions and methods for personalization. The effectiveness of our adversarial watermark defense through personalized fine-tuning is observed across different model versions and personalized methods.
Main Results. For each image processing attack, we report the average bit accuracy/ TPR@1e-5FPR in Table 2, which shows ConceptWM is indeed robust across all the transformations with the bit accuracy consistently above 0.9. ConceptWM remarkably outperforms existing multi-bit watermarks on the average performance with more than 6.3% boost compared with the state-of-the-art (SoTA) results. Since our method embeds the watermark into the entire latent space and trains the distortion layer, it has a clear advantage over contrastive watermarking schemes when facing various image processing attacks.
@@
Table 5. The impact of various attacks on image quality in the distortion layer in the watermark components pretraining.
4.5. Concept Protection Evaluation
@@
Textual Inversion and DreamBooth with LoRA. Textual Inversion teaches new concepts by adjusting a word vector, rather than fine-tuning the entire model. LoRA, on the other hand, employs low-rank weight updates to enhance memory efficiency and is widely adopted within the community. The relevant results are shown in Table 4. Our method shows better performance in mitigating the degradation effect on generation with LoRA. Since Textual Inversion does not inject concepts directly into the model, the reduction in quality is not significant and the semantic quality of the generation decrease.
@@
Inference Steps. Users may influence watermark extraction under different inference parameters. We generated images of specific concepts at different inference steps and performed watermark extraction. The results, shown in Table 3, indicate that the ConceptWM remains effective even with varying inference steps.
@@
Sampling Method. We evaluate our watermark on DDIM (Song et al., 2020), DPM-Solver multi-step (Lu et al., 2022), Euler and Heun samplers (Karras et al., 2022). The results show that our watermark is minimally affected by the choice of sampler.
@@
Generation Size. The results in Table 3 show that different generation sizes have a noticeable impact on watermark, while the watermarking method still remains effective and detectable across different generation sizes.
5. Conclusion
This paper reveals the potential threats posed by diffusion model personalization techniques and proposes an adversarial watermarking method to protect specific concepts. Our solution integrates adversarial perturbations and watermarks into the diffusion model, ensuring that images generated by the diffusion model for specific concepts are traceable, and any personalized model trained on these images will produce low-quality results. We have evaluated the stability and adversarial effectiveness of the watermarking method we designed, demonstrating that our defense remains effective under adverse conditions. In the future, our main focus will be on two aspects of improvement. On the one hand, we aim to address the efficiency issue caused by the need to generate adversarial noise in the current method. On the other hand, we will consider potential threats from attackers who may target the model level.